Carlos is three times as old as Maria. The sum of their ages is 56 years. How old is each?
Answer by JulietG(1812)

C = 3M
C + M = 56
Substitute the value of C from the first equation into the second.
C + M = 56
(3M) + M = 56
4M = 56
Divide each side by 4
M = 14
If Maria is 14, then Carlos is 42